A 15 A current flows in a device that has a resistance of

0.15Ω. Compute the rate of heat production by the device.


1
Expert's answer
2022-05-30T15:18:24-0400

The rate of heat production by the device

"P=I^2R\\
=\rm (15\: A)^2\times 0.15\:\Omega=34\: W"